# NeuralSurv: Deep Survival Analysis with Bayesian Uncertainty Quantification

### v1

@@ Abstract
We introduce NeuralSurv, the first deep survival model to incorporate Bayesian uncertainty quantification. Our non-parametric, architecture-agnostic framework flexibly captures time-varying covariate–risk relationships in continuous time via a novel two-stage data-augmentation scheme, for which we establish theoretical guarantees. For efficient posterior inference, we introduce a mean-field variational algorithm with coordinate-ascent updates that scale linearly in model size. By locally linearizing the Bayesian neural network, we obtain full conjugacy and derive all coordinate updates in closed form. In experiments, NeuralSurv delivers superior calibration compared to state-of-the-art deep survival models, while matching or exceeding their discriminative performance across both synthetic benchmarks and real-world datasets. Our results demonstrate the value of Bayesian principles in data-scarce regimes by enhancing model calibration and providing robust, well-calibrated uncertainty estimates for the survival function.
@@ 1 Introduction
Survival analysis is a branch of statistics focused on the study of time-to-event data, usually called event times. This type of data appears in a wide range of applications such as medicine survival_analysis_medicine_book , engineering survival_book_engineering , and social sciences survival_analysis_social_sciences . A key objective of survival analysis is to estimate the survival function and the hazard function that govern the distribution of event times.
@@
Traditional survival models like the Cox proportional hazards model Cox1972 and accelerated failure time models Carroll2003 have long delivered reliable inference under strong parametric assumptions. However, such assumptions may fail to adequately capture complex and evolving baseline hazards, especially when risk relationships vary over time. To overcome these limitations, recent work has begun incorporating modern machine-learning techniques Wang2019 , and in particular deep architectures Wiegrebe2024 ; Katzman2018 ; Lee2018 , which can learn rich, hierarchical representations directly from data. Yet most deep-survival approaches remain purely frequentist, optimizing point-estimate losses and offering no coherent uncertainty quantification. In high-stakes settings like medicine, this lack of reliable uncertainty estimates can undermine trust and impede adoption.
@@
Bayesian statistics, by contrast, inherently quantifies uncertainty: prior beliefs are combined with observed data to yield a posterior distribution over model parameters bda_book . In survival analysis, Bayesian methods can produce full posterior distributions for individual survival functions summarizable via credible intervals that communicate model confidence bayesian_survival_analysis . Traditional Bayesian survival tools, such as Gaussian processes (GPs) teh_survival ; Kim2018 , offer nonparametric flexibility and built-in uncertainty but often falter in high-dimensional settings due to scalability issues. To date, no method has married the representational power of deep learning with full Bayesian uncertainty quantification in a scalable survival framework. Such a synthesis would hold the potential to learn complex, high-dimensional survival dynamics while retaining principled probabilistic interpretations.
@@
In this work, we address this critical need by introducing NeuralSurv, an architecture-agnostic, Bayesian deep-learning framework for survival analysis. NeuralSurv leverages deep neural networks to learn hierarchical representations from covariates and uses a principled variational inference framework to provide rigorous uncertainty quantification over the survival function. We develop a novel two-stage data-augmentation strategy that leverages latent marked Poisson processes and Pólya–Gamma variables. Our approach comes with theoretical guarantees and enables exact continuous-time likelihood computation. By locally linearizing the Bayesian neural network, we achieve conjugacy and derive closed-form coordinate-ascent updates that scale linearly with network size.
@@
Through extensive experiments on synthetic and real survival datasets, in data-scarce settings, NeuralSurv consistently delivers superior calibration compared to state-of-the-art deep survival models, and matches or exceeds their discriminative performance. Its Bayesian formulation captures epistemic uncertainty to prevent overfitting, while informative priors induce a soft regularization that yields smooth, plausible survival functions.
@@ 2 NeuralSurv
We briefly introduce our survival analysis setup; for a comprehensive review, see Appendix A. Let $(T_{i},C_{i})$ denote the survival and censoring times respectively for observations $i=1,\ldots,N$. We observe $\mathcal{D}=\{(y_{i},\delta_{i}):i=1,\ldots,N\}$, where $y_{i}=\min(T_{i},C_{i})$ is the observed (right-censored) event time and $\delta_{i}=\mathbbm{1}_{\{T_{i}\leq C_{i}\}}$ is the censoring indicator. Each observation has a covariate vector $\mathbf{x}_{i}\in\mathbb{R}^{p}$, collected into $\mathbf{X}=\{\mathbf{x}_{i}:i=1,\ldots,N\}$. Throughout this paper, we assume that the censoring time $C_{i}$ is independent of the event time $T_{i}$ given $\mathbf{x}_{i}$.
Our goal is to model the hazard function $\lambda$, i.e. the instantaneous event rate at time $t$ conditional on survival to $t$ and covariates $\mathbf{x}$. We assume a sigmoidal hazard function:
$\lambda(t\mid\mathbf{x};\phi,g(\cdot;\boldsymbol{\theta}))=\lambda_{0}(t,\mathbf{x};\phi)\>\sigma(g(t,\mathbf{x};\boldsymbol{\theta})),$ (1)
where the sigmoid function $\sigma(z)=1/(1+\exp(-z))$ maps any real input to a value in $(0,1)$. Here, the baseline hazard $\lambda_{0}:\mathbb{R}_{+}\times\mathbb{R}^{p}\to\mathbb{R}_{+}$, parametrized by ϕ+italic-$\phi\in\mathbb{R}_{+}$, encodes our prior “best-guess” hazard profile over time and covariates. Finally, the flexible function $g:\mathbb{R}_{+}\times\mathbb{R}^{p}\to\mathbb{R}$, parametrized by $\boldsymbol{\theta}\in\mathbb{R}^{m}$, provides a data-driven adjustment: once passed through the sigmoid, it multiplicatively modulates the baseline hazard, continuously scaling it between zero and $\lambda_{0}$.
@@ 2.1 Likelihood Distribution
Assuming $(y_{i},\delta_{i})$ are i.i.d. conditional on $(\mathbf{x}_{i},\phi,g(\cdot;\boldsymbol{\theta}))$, the full-sample likelihood is simply the product
@@ 2.2 Prior Distributions
We assume that $g(\cdot;{\boldsymbol{\theta}})$ is a Bayesian Neural Network (BNN) parameterized by $\boldsymbol{\theta}$. Furthermore, denote by $\mathbf{I}_{m}$ the $m\times m$ identity matrix. We place the following isotropic Gaussian prior with zero mean and identity covariance over the network weights
@@
We further assume that the baseline hazard can be written as
$\lambda_{0}(t,\mathbf{x};\phi)=\frac{\lambda_{0}(t;\phi)}{Z(t,\mathbf{x})},$ (5)
for the normalization function
$Z(t,\mathbf{x}):=\mathbb{E}_{\boldsymbol{\theta}\sim p_{\boldsymbol{\theta}}}\left[\sigma(g(t,\mathbf{x};\boldsymbol{\theta}))\right];$
see Appendix D for further details on how to obtain $Z(t,\mathbf{x})$. By introducing the normalizing function $Z(t,\mathbf{x})$ in (5) we ensure that the prior mean of the sigmoidal hazard in (1) coincides with the prior mean of the baseline hazard in (5), i.e.
@@
This approach, similar to the technique used in teh_survival , centers the distribution around the baseline hazard $\lambda_{0}(t;\phi)$, favouring hazard trajectories that remain close to this prior “best-guess” profile while still permitting data-driven deviations. Notice that if $g(\cdot)$ has a fully connected architecture, then $Z(t,\mathbf{x})\equiv\frac{1}{2}$ for all $(t,\mathbf{x})$, resulting in the same normalization function value as in teh_survival .
We adopt a Weibull-type baseline hazard
where $\alpha_{0}$ is the shape and $\beta_{0}$ is the rate of the Gamma distribution. The Weibull-type baseline hazard (6) is the hazard of a Weibull distribution, a common choice in survival analysis teh_survival . When $\rho=1$, $\lambda_{0}(t;\phi)$ becomes constant and the baseline hazard reduces to the hazard of the Exponential distribution.
@@ 2.3 Posterior Distribution
The posterior in (7) is generally intractable to compute for three reasons. First, its normalizing constant is unavailable in closed form. Second, the likelihood from (2) requires evaluating $N$ integrals, none of which admits an analytic solution. Finally, the sigmoid in (1) introduces an extra nonlinearity, rendering inference even more analytically challenging.
@@ 3 Data Augmentation Strategy
In this section, we present a data augmentation scheme that leverages the properties of Poisson processes and Pólya-Gamma random variables. Specifically, Poisson processes help overcome the challenges associated with computing the integral of the continuous-time function to evaluate the likelihood, while the Pólya-Gamma random variables allow for exact handling of the sigmoid nonlinearity without relying on analytic approximations. This combined approach allows us to efficiently perform posterior inference from the model without resorting to discretization. In Section 4.3, we develop a novel variational inference algorithm based on this augmentation scheme. Detailed reviews of Pólya-Gamma random variables and Poisson processes are provided in Appendices B and C, respectively.
@@ 3.2 Poisson Process Augmentation Scheme
Evaluating the likelihood in (2) requires computing an integral involving a sample function drawn from the BNN prior. This integral is generally analytically intractable, due to the nonparametric and highly non-linear nature of BNN sample paths. To address this, we propose a Poisson process augmentation scheme. By substituting the sigmoid identity from (9), the intractable integral for the $i^{\text{th}}$ data point becomes
$\int_{0}^{y_{i}}\lambda_{0}(t,\mathbf{x}_{i};\phi)\>\sigma(g(t,\mathbf{x}_{i};\boldsymbol{\theta}))\mathrm{d}t=\\ \int_{0}^{y_{i}}\int_{0}^{\infty}\left(1-e^{f(\omega,-g(t,\mathbf{x}_{i};\boldsymbol{\theta}))}\right)\lambda_{0}(t,\mathbf{x}_{i};\phi)p_{\text{PG}}(\omega\mid 1,0)\mathrm{d}\omega\mathrm{d}t,$
@@
Before proceeding further, we briefly review the concept of a marked Poisson process. A marked Poisson process extends the standard Poisson process by associating each event (or location) with an additional random variable known as a mark. In our case, each event occurs at time $t$ and is accompanied by a positive mark $\omega$. With this in mind, consider the space $[0,y_{i}]\times\mathbb{R}_{+}$ which consists of points $(t,\omega)$ where $t\in[0,y_{i}]$ and $\omega\in\mathbb{R}_{+}$. We then denote by $\Psi_{i}$ a marked Poisson process on $[0,y_{i}]\times\mathbb{R}_{+}$ with intensity
@@
Under suitable assumptions on the BNN $g(t,\mathbf{x}_{i};\boldsymbol{\theta})$, Campbell’s theorem allows us to express the integral as
@@ 3.3 Augmented Likelihood
The proof of Theorem 3.1 is postponed to Appendix N.1. Analogous data augmentation schemes have been proposed in donner_opper_gaussian_cox ; pg_poisson_hawkes ; mutually_regressive_pp . However, to the best of our knowledge, this is the first application of such an approach to survival analysis, and the first to provide a rigorous theoretical framework that establishes the validity of the methodology.
@@ 4.1 Variational Mean–Field Approximation
Computing the posterior distribution $\mathbb{P}\left(\phi,\boldsymbol{\theta},\boldsymbol{\omega},\boldsymbol{\Psi}\mid\mathcal{D},\mathbf{X}\right)$ is analytically intractable. Therefore, we consider a variational inference algorithm that aims to find an approximating variational distribution $\mathbb{Q}(\phi,\boldsymbol{\theta},\boldsymbol{\omega},\boldsymbol{\Psi})$ that minimizes the KL divergence from the true posterior distribution.
@@
and where $\frac{\mathrm{d}\mathbb{P}_{\boldsymbol{\Psi}|\phi}}{\mathrm{d}\mathbb{P}_{\boldsymbol{\Psi},*}}$ is the Radon-Nykodim derivative of the true conditional law $\mathbb{P}_{\boldsymbol{\Psi}|\phi}$ with respect to $\mathbb{P}_{\boldsymbol{\Psi},*}$, cf.  (28). From (15), it follows that minimizing the KL divergence is equivalent to maximizing the ELBO.

### v2

@@ Abstract
We introduce NeuralSurv, the first deep survival model to incorporate Bayesian uncertainty quantification. Our non‑parametric, architecture‑agnostic framework captures time‑varying covariate–risk relationships in continuous time via a novel two‑stage data‑augmentation scheme, for which we establish theoretical guarantees. For efficient posterior inference, we introduce a mean‑field variational algorithm with coordinate‑ascent updates that scale linearly in model size. By locally linearizing the Bayesian neural network, we obtain full conjugacy and derive all coordinate updates in closed form. In experiments, NeuralSurv delivers superior calibration compared to state-of-the-art deep survival models, while matching or exceeding their discriminative performance across both synthetic benchmarks and real-world datasets. Our results demonstrate the value of Bayesian principles in data‑scarce regimes by enhancing model calibration and providing robust, well‑calibrated uncertainty estimates for the survival function.
@@ 1 Introduction
Survival analysis is a branch of statistics focused on the study of time-to-event data, usually called event times. This type of data appears in a wide range of applications such as medicine survival_analysis_medicine_book, engineering survival_book_engineering, and social sciences survival_analysis_social_sciences. A key objective of survival analysis is to estimate the hazard function and the survival function that govern the distribution of event times.
@@
Traditional survival models like the Cox proportional hazards model Cox1972 and accelerated failure time models Carroll2003 have long delivered reliable inference under strong parametric assumptions. However, such assumptions may fail to adequately capture complex and evolving baseline hazards, especially when risk relationships vary over time. To overcome these limitations, recent work has begun incorporating modern machine‑learning techniques Wang2019, and in particular deep architectures Wiegrebe2024; Katzman2018; Lee2018, which can learn rich, hierarchical representations directly from data. Yet most deep‑survival approaches remain purely frequentist, optimizing point‑estimate losses and offering no coherent uncertainty quantification. In high‑stakes settings like medicine, this lack of reliable uncertainty estimates can undermine trust and impede adoption.
@@
Bayesian statistics, by contrast, inherently quantifies uncertainty: prior beliefs are combined with observed data to yield a posterior distribution over model parameters bda_book. In survival analysis, Bayesian methods can produce full posterior distributions for individual survival functions summarizable via credible intervals that communicate model confidence bayesian_survival_analysis. Traditional Bayesian survival tools, such as Gaussian processes (GPs) teh_survival; Kim2018, offer nonparametric flexibility and built‑in uncertainty but often falter in high‑dimensional settings due to scalability issues. To date, no method has combined the representational power of deep learning with full Bayesian uncertainty quantification in a scalable survival framework. Such a synthesis would hold the potential to learn complex, high‑dimensional survival dynamics while retaining principled probabilistic interpretations.
@@
In this work, we introduce NeuralSurv, an architecture‑agnostic, Bayesian deep‑learning framework for survival analysis which integrates with modern deep learning architectures. NeuralSurv leverages deep Neural Networks (NNs) to learn hierarchical representations from covariates and uses a principled variational inference framework to provide rigorous uncertainty quantification over the survival function. We develop a two-stage data-augmentation strategy using latent marked Poisson processes and Pólya–Gamma variables to enable exact continuous-time likelihood computation, and provide novel theoretical guarantees for this approach. By locally linearizing the Bayesian Neural Network (BNN), we achieve conjugacy and derive closed‑form coordinate‑ascent updates that scale linearly with network size.
@@
Through extensive experiments on synthetic and real survival datasets, in data‐scarce settings, NeuralSurv consistently delivers superior calibration compared to state‑of‑the‑art deep survival models, and matches or exceeds their discriminative performance. Its Bayesian formulation captures epistemic uncertainty to prevent overfitting, while informative priors induce a soft regularization that yields smooth, plausible survival functions. The code to reproduce our experiments is available on the GitHub repository https://github.com/MLGlobalHealth/neuralsurv under the MIT License.
@@ 2 NeuralSurv
In this section, we outline the main assumptions underlying NeuralSurv. We begin by briefly reviewing key concepts in survival analysis. Survival analysis focuses on modeling time-to-event data. Let $T$ be a continuous nonnegative random variable with probability density $f$ and cumulative distribution function $F$, representing the time until a particular event occurs. Its survival function $S(t)=\mathbb{P}(T>t)=1-F(t)$ gives the probability of not experiencing the event by time $t$, while the hazard function $\lambda(t)=f(t)/S(t)$ represents the instantaneous risk of the event at time $t$, conditional on having survived up to time $t$. In practice, the event time may not be observed for all individuals, because some observations are subject to right-censoring, where the event has not yet occurred by the end of the observation period. For each observation $i=1,\ldots,N$, denote the event time by $T_{i}$ and the censoring time by $C_{i}$. We observe $y_{i}=\min(T_{i},C_{i})$ and $\delta_{i}=\mathbbm{1}_{\{T_{i}\leq C_{i}\}}$ where $y_{i}$ represents the observed time (which may correspond either to the event or to censoring), and $\delta_{i}$ indicates whether the event time was observed ($\delta_{i}=1$) or the observation period was censored ($\delta_{i}=0$). We assemble the dataset as $\mathcal{D}=\{(y_{i},\delta_{i}):i=1,\ldots,N\}$. Each observation also carries a covariate vector $\mathbf{x}_{i}\in\mathbb{R}^{p}$, gathered into $\mathbf{X}=\{\mathbf{x}_{i}:i=1,\ldots,N\}$. Throughout this paper, we assume that the censoring time $C_{i}$ is independent of the event time $T_{i}$ given $\mathbf{x}_{i}$ (known as non-informative censoring). Further details on survival analysis theory are provided in Appendix A.
2.1 Sigmoidal Hazard Function
Our goal is to model the hazard function $\lambda$, i.e. the instantaneous event rate at time $t$ conditional on survival to $t$ and covariates $\mathbf{x}$. We employ the sigmoid function $\sigma(z)=1/(1+\exp(-z))$, which maps real-valued inputs to the interval $(0,1)$. The sigmoidal hazard model is constructed as the product of a normalized baseline hazard function ($\lambda_{0}$) and a modulation function ($\sigma$):
$\lambda(t\mid\mathbf{x};\phi,g(\cdot;\boldsymbol{\theta})):=\lambda_{0}(t,\mathbf{x};\phi)\ \sigma(g(t,\mathbf{x};\boldsymbol{\theta})),$ (1)
where the normalized baseline hazard is given by
$\lambda_{0}(t,\mathbf{x};\phi):=\frac{\lambda_{0}(t;\phi)}{Z(t,\mathbf{x})},$ (2)
for the baseline hazard $\lambda_{0}:\mathbb{R}_{+}\to\mathbb{R}_{+}$, parametrized by $\phi\in\mathbb{R}_{+}$, and a normalization factor $Z(t,\mathbf{x})$ that depends on both time and covariates. The term $\lambda_{0}(t,\mathbf{x};\phi)$ encodes our prior “best‐guess” hazard profile over time. The flexible function $g:\mathbb{R}_{+}\times\mathbb{R}^{p}\to\mathbb{R}$, parametrized by $\boldsymbol{\theta}\in\mathbb{R}^{m}$, provides a data‐driven adjustment: once passed through the sigmoid, it multiplicatively attenuates the baseline hazard, continuously scaling it between zero and $\lambda_{0}$. The normalization factor $Z(t,\mathbf{x})$ ensures that the overall hazard remains properly scaled after modulation by the sigmoidal function (see Section 2.3 for details). Modeling hazard and intensity functions using sigmoidal transformations is common in applications such as survival analysis teh_survival and point process models pg_poisson_hawkes; donner_opper_gaussian_cox; renewal_process; Adams2009; aglietti_cox. This approach is popular due to the balance it offers between modeling flexibility and analytical tractability.
@@ 2.2 Likelihood Distribution
Assuming $(y_{i},\delta_{i})$ are i.i.d. conditional on $(\mathbf{x}_{i},\phi,g(\cdot;\boldsymbol{\theta}))$, the full‐sample likelihood is simply the product
@@ 2.3 Prior Distributions
We assume that $g(\cdot;{\boldsymbol{\theta}})$ is a BNN parameterized by $\boldsymbol{\theta}$. Furthermore, denote by $\mathbf{I}_{m}$ the $m\times m$ identity matrix. We place the following isotropic Gaussian prior with zero mean and identity covariance over the NN weights
@@
We adopt a Weibull‑type baseline hazard
where $\alpha_{0}$ is the shape and $\beta_{0}$ is the rate of the Gamma distribution. The Weibull‑type baseline hazard (6) is the hazard of a Weibull distribution, a common choice in survival analysis teh_survival. When $\rho=1$, $\lambda_{0}(t;\phi)$ becomes constant and the baseline hazard reduces to the hazard of the Exponential distribution.
Normalization Factor $Z$.
We define the normalization factor introduced in (2) as
$Z(t,\mathbf{x}):=\mathbb{E}_{\boldsymbol{\theta}\sim p_{\boldsymbol{\theta}}}\left[\sigma(g(t,\mathbf{x};\boldsymbol{\theta}))\right]$
and refer the reader to Appendix D for further details on how it is computed. Introducing this normalization factor ensures that the prior mean of the sigmoidal hazard in (1) coincides with the prior mean of the baseline hazard, i.e.
@@
This approach, similar to the technique used in teh_survival, centers the distribution around the baseline hazard $\lambda_{0}(t;\phi)$, favouring hazard trajectories that remain close to this prior “best‐guess” profile while still permitting data‐driven deviations. Notice that if $g(\cdot;\boldsymbol{\theta})$ has a fully connected architecture, then $Z(t,\mathbf{x})\equiv\frac{1}{2}$ for all $(t,\mathbf{x})$, resulting in the same normalization factor value as in teh_survival.
@@ 2.4 Posterior Distribution
The posterior in (7) is generally intractable to compute for three reasons. First, its normalization constant is unavailable in closed form. Second, the likelihood from (3) requires evaluating $N$ integrals, none of which admits an analytic solution. Finally, the sigmoid in (1) introduces an extra nonlinearity, rendering inference even more analytically challenging.
@@ 3 Data Augmentation Strategy
In this section, we present a data augmentation scheme that leverages the properties of Poisson processes and Pólya-Gamma random variables. Specifically, Poisson processes help overcome the challenges associated with computing the integrals of the continuous-time function to evaluate the likelihood, while the Pólya-Gamma random variables allow for exact handling of the sigmoid nonlinearity without relying on analytic approximations. This combined approach allows us to efficiently perform posterior inference from the model without resorting to discretization.
This approach builds on analogous strategies previously applied in other settings, including Bayesian inference for Sigmoid Gaussian Cox Processes donner_opper_gaussian_cox, nonparametric Hawkes processes pg_poisson_hawkes, and, in the case of Pólya–Gamma augmentation alone, mutually regressive point processes mutually_regressive_pp. To the best of our knowledge, this is the first application of such a data augmentation strategy in the context of survival analysis. Furthermore, we are the first to provide a rigorous theoretical framework that establishes the validity of a method belonging to this broader class of augmentation-based approaches (see Theorem 3.1).
Detailed reviews of Pólya-Gamma random variables and Poisson processes are provided in Appendices B and C, respectively.
@@ 3.2 Poisson Process Augmentation Scheme
Evaluating the likelihood in (3) requires computing $N$ integrals involving a sample function drawn from the BNN prior. This integral is generally analytically intractable, due to the nonparametric and highly non-linear nature of BNN sample paths. To address this, we leverage a Poisson process–based data augmentation scheme, drawing inspiration from methodologies proposed in donner_opper_gaussian_cox; pg_poisson_hawkes. By substituting the sigmoid identity from (9), the intractable integral for the $i^{\text{th}}$ data point becomes
$\int_{0}^{y_{i}}\lambda_{0}(t,\mathbf{x}_{i};\phi)\>\sigma(g(t,\mathbf{x}_{i};\boldsymbol{\theta}))\mathrm{d}t=\\ \int_{0}^{y_{i}}\int_{0}^{\infty}\left(1-e^{f(\omega,-g(t,\mathbf{x}_{i};\boldsymbol{\theta}))}\right)\lambda_{0}(t,\mathbf{x}_{i};\phi)p_{\text{PG}}(\omega\mid 1,0)\mathrm{d}\omega\mathrm{d}t,$ (11)
@@
Before proceeding further, we briefly review the concept of a marked Poisson process. A marked Poisson process extends the standard Poisson process by associating each event (or location) with an additional random variable known as a mark. In our case, each event occurs at time $t$ and is accompanied by a positive mark $\omega$. With this in mind, consider the space $0,y_{i}]\times\mathbb{R}_{+}$ which consists of points $(t,\omega)$ where $t\in[0,y_{i}]$ and $\omega\in\mathbb{R}_{+}$. We then denote by $\Psi_{i}$ a marked Poisson process on $0,y_{i}]\times\mathbb{R}_{+}$ with intensity
@@
Under suitable assumptions on the BNN $g(\cdot;\boldsymbol{\theta})$, Campbell’s theorem allows us to express the integral in (11) as
@@ 3.3 Augmented Likelihood
The proof of Theorem 3.1 is postponed to Appendix N.1. Existing augmentation schemes approaches donner_opper_gaussian_cox; pg_poisson_hawkes; mutually_regressive_pp do not offer any theoretical guarantees regarding the validity of the methodology. In contrast, Theorem 3.1 provides the first rigorous theoretical framework that establishes the soundness of a method within this class of data augmentation techniques.
@@ 4 Variational Inference in the Augmented Space
In this section, we develop a novel variational inference algorithm based on this augmentation scheme.
@@ 4.1 Variational Mean–Field Approximation
Computing the posterior distribution $\mathbb{P}\left(\phi,\boldsymbol{\theta},\boldsymbol{\omega},\boldsymbol{\Psi}\mid\mathcal{D},\mathbf{X}\right)$ is analytically intractable because its normalization constant is unavailable in closed form. We consider a variational inference algorithm that aims to find an approximating variational distribution $\mathbb{Q}(\phi,\boldsymbol{\theta},\boldsymbol{\omega},\boldsymbol{\Psi})$ that minimizes the KL divergence from the true posterior distribution.
@@
and where $\frac{\mathrm{d}\mathbb{P}_{\boldsymbol{\Psi}|\phi}}{\mathrm{d}\mathbb{P}_{\boldsymbol{\Psi},*}}$ is the Radon-Nykodim derivative of the true conditional law $\mathbb{P}_{\boldsymbol{\Psi}|\phi}$ with respect to $\mathbb{P}_{\boldsymbol{\Psi},*}$. From (16), it follows that minimizing the KL divergence is equivalent to maximizing the ELBO.